OpenStudy (anonymous):

Please help me????????????? Solve by using the quadratic formula to solve x2 + 6x = -3.

OpenStudy (anonymous):

easier to complete the square \[x^2+6x=-3\] \[(x+3)^2=-3+9=6\] \[x+3=\pm\sqrt{6}\] \[x=-3\pm\sqrt{6}\]

OpenStudy (anonymous):

x=-0.5505 , x=-5.449

OpenStudy (anonymous):

write it like this \[x^{2} + 6x +3=0\] a=1 b=6 c=3, now you cam apply the quadretic formula \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

omg thank you guys........!!!!!!!!!!!!!!!!!!!!!!!!!!!!!